# Separate Motion from Appearance: Customizing Motion via Customizing Text-to-Video Diffusion Models

### v1

@@ Abstract
Motion customization aims to adapt the diffusion model (DM) to generate videos with the motion specified by a set of video clips with the same motion concept. To realize this goal, the adaptation of DM should be possible to model the specified motion concept, without compromising the ability to generate diverse appearances. Thus, the key to solving this problem lies in how to separate the motion concept from the appearance in the adaptation process of DM. Typical previous works explore different ways to represent and insert a motion concept into large-scale pretrained text-to-video diffusion models, e.g., learning a motion LoRA, using latent noise residuals, etc. While those methods can encode the motion concept, they also inevitably encode the appearance in the reference videos, resulting in weakened appearance generation capability. In this paper, we follow the typical way to learn a motion LoRA to encode the motion concept, but propose two novel strategies to enhance motion-appearance separation, including temporal attention purification (TAP) and appearance highway (AH). Specifically, we assume that in the temporal attention module, the pretrained Value embeddings are sufficient to serve as basic components needed by producing a new motion. Thus, in TAP, we choose only to reshape the temporal attention with motion LoRAs so that Value embeddings can be reorganized to produce a new motion. Further, in AH, we alter the starting point of each skip connection in U-Net from the output of each temporal attention module to the output of each spatial attention module. Extensive experiments demonstrate that compared to previous works, our method can generate videos with appearance more aligned with the text descriptions and motion more consistent with the reference videos.
@@
Figure 1: Limitations of previous motion customization methods. Previous methods [47] suffer from the issue of “appearance leakage”, where elements from the reference videos, such as “window”, unexpectedly appear in the generated video.
1 Introduction
@@
Text-to-video diffusion models [5, 3] have made remarkable progress. With just a text input, these models can generate high-quality videos that faithfully reflect the prompt provided. However, relying solely on text descriptions, pretrained text-to-video diffusion models struggle to generate high-fidelity videos for highly customized motion concepts. To address this issue, researchers explore various ways to realize motion customization of text-to-video diffusion models.
@@
Formally, motion customization aims to adapt the diffusion model (DM) to generate videos with the motion specified by a set of video clips (i.e., reference videos) with the same motion concept. To realize this goal, the adaptation of DM should be possible to model the specified motion concept, without compromising the ability to generate diverse appearance. Thus, the key to solving this problem lies in how to separate the motion concept from the appearance in the adaptation process of diffusion models. Typical previous works [47, 24, 18] explore different ways to represent and insert a motion concept into large-scale pretrained text-to-video diffusion models. For example,  [47] trains motion LoRAs to encode the motion concept from the reference videos. Additionally, [47] trains appearance LoRAs to encode the appearance of reference videos. At inference, the appearance LoRAs are abandoned to avoid overfitting to the appearance of reference videos. Other works like [18] utilize residual vectors between consecutive frames to represent a motion. While those methods can encode the motion concept, there exists no explicit constraint that appearance information should not be learned or encoded into the motion representations. Actually, they inevitably encode the appearance of the reference videos (termed as “appearance leakage”), resulting in weakened appearance generation ability. As shown in the Fig. 1 (first row), the background of the reference videos includes some “window”, which has nothing to do with the motion concept. Previous methods (second row), while fitting the motion, also mistakenly fit the appearance of the “window”. As a result, the generated video’s background also includes “window”, which is not aligned with the text description.
@@
In this paper, we follow the typical way to learn motion LoRAs to encode the motion concept but propose two novel strategies to enhance motion-appearance separation, including temporal attention purification (TAP) and appearance highway (AH). We choose to adopt the typical text-to-video diffusion model (T2V-DM) ZeroScope [31] or ModelScope [34] to perform the motion customization. Specifically, we assume that in the temporal transformers of text-to-video diffusion models, the pretrained Value embeddings are sufficient to serve as basic components needed to produce a new motion. Thus, in TAP, we choose only to reshape the temporal attention with motion LoRAs so that Value embeddings can be reorganized to produce a new motion. As TAP may not perfectly avoid encoding appearance information in motion LoRAs, we propose AH to maintain the appearance generation ability of original T2V-DM and further avoid appearance leakage issue. Specifically, we alter the starting point of each skip connection in spatial-temporal U-Net [8, 14] from the output of each temporal transformers to the output of each spatial transformers. The underlying logic of AH is that as the skip connection of U-Net mainly conveys high-frequency appearance information [28], the starting point alternation can provide a “highway” for the hidden states of non-adapted spatial transformers and cut the shortcut from the hidden states of LoRA-adapted temporal transformers. We empirically find that through AH, the hidden states from the decoding branch of spatial-temporal U-Net can encode appearance information more similar to the vanilla T2V-DM and motion information more similar to the T2V-DM adapted with TAP. This means AH largely maintains the appearance generation ability of vanilla T2V-DM, without harming the motion concept learning from reference videos. With the proposed TAP and AH, our method can largely mitigate the appearance leakage issue and generate videos with customized motion and high-fidelity appearance well aligned with the text descriptions, as shown in Fig. 1 (last row). Additionally, we point out that motion modeling occurs in the early steps of the denoising process and propose phased LoRA integration (PIL) to maximize the T2V-DM’s appearance modeling capability. Extensive experiments demonstrate that our method effectively separates motion from appearance and outperforms previous state-of-the-arts.
In a nutshell, our contributions are summarized as
@@
  • We conducted extensive qualitative and quantitative experiments, demonstrating that our method outperforms previous state-of-the-arts.
@@
Figure 2: Overview of our method. This figure illustrates our main contributions: (a) Overall inference process. (b) Temporal Attention Purification where we utilize LoRA only to adapt the Key embeddings to reshape the temporal attention. (c) Appearance Highway where we alter the starting point of skip connection from LoRA-adapted temporal transformer block to non-adapted spatial transformer block.
2 Related Work
2.1 Text-to-Video generation.
@@ 2 Related Work
Text-to-video generation is a task that involves generating realistic videos based on text prompts. Early research primarily focused on Generative Adversarial Networks (GAN) [33, 2, 32, 27], autoregression transformers[42, 15, 21, 10]. Recently, diffusion models have emerged as the most popular paradigm in text-to-video generation [5, 1, 9, 16, 44, 45, 46, 3]. Make-A-Video [29] is first trained on labeled images and then on unlabeled videos to address the issue of the lack of paired video-text data. Imagen Video[13] generates high-quality videos by using a text-to-video model combined with spatial and temporal super-resolution models. MagicVideo[48] and VideoLDM[4] adapts the Latent Diffusion Model’s [26] architecture for video generation. In contrast to these methods, Text2Video-Zero[19], AnimateDiff[12] leverage the prior knowledge of image generation models to generate videos. Recently, open-source T2V-DM, ModelScope [34] and ZeroScope [31] has attracted considerable attention.
2.2 Motion customization generation.
@@
To generate motion-customized videos, several video editing methods [6, 43, 11, 22], including Control-A-Video [7] and VideoComposer [35], transfer motion from a reference video to new appearances. These approaches primarily modify the appearance without fully capturing the essence of motion, which limits appearance diversity. Tune-A-Video [38] extends text-to-image models into the video domain, enabling customized video generation. VMC [18] utilizes noise residuals to represent motion information and minimizes residual loss to fine-tune pretrained text-to-video models. LAMP [40] and DreamVideo [37] incorporate an additional image paired with editing text as input, serving as guidance for the generation process. However, this approach increases the complexity of user input requirements. To address these challenges, MotionDirector [47] and Customize-a-Video [24] introduce prior knowledge of video appearance into their models, encouraging a stronger focus on learning motion rather than appearance. Despite these advancements, achieving a complete disentanglement of motion and appearance remains problematic, with models often overfitting to the reference video’s appearance.
3 Method
@@
In this paper, we aim to adapt the text-to-video diffusion model (T2V-DM) to perform motion customization, i.e., we need to adapt T2V-DM with the reference videos containing one specific motion concept to generate new high-fidelity videos with the same motion concept. The general framework of our method is illustrated in Fig. 2. Our method is based on typical pre-trained T2V DMs (e.g., ZeroScope [31] or ModelScope [34]) whose architecture is a 3D U-Net. The basic components for such a 3D U-Net include spatial transformer and temporal transformer, which encode the spatial and temporal information respectively. Skip connections are utilized to connect the encoders and decoders of the U-Net.
@@
Following MotionDirector, we adopt a dual-path LoRA adaption for training (Sec. 3.1), which tunes the model with low-rank adaptions (LoRAs) while keeping the pre-trained parameters fixed, and finally models the motion information in the LoRAs of temporal transformers. However, such a strategy still inevitably encodes the appearance of the reference videos into temporal LoRAs, resulting in weakened appearance generation ability. Thus, we propose two effective modifications. Firstly, instead of injecting LoRAs into all linear layers of the attention mechanism in temporal transformers, we propose a Temporal Attention Purification (TAP) strategy (Sec. 3.2). Secondly, at the inference stage, we modify the skip connection between the encoder and decoder of spatial-temporal U-Net with an Appearance Highway strategy (Sec. 3.3), which transfers the hidden states from the spatial transformers rather than the temporal transformers of the encoder to the decoder. Finally, during inference, we adopt a phased LoRA integration trick where we utilize our modified framework to denoise for a few steps and then use the original spatial-temporal U-Net for the latter steps of denoising (Sec. 2).
3.1 Baseline: Dual-Path LoRA Adaption
@@
MotionDirector [47] utilizes a spatial-temporal U-Net [8, 14], which takes visual latent code $z_{t}\in\mathcal{R}^{b\times f\times h\times w\times c}$ ($b$, $f$, $h$, $w$, $c$ represents batch size, frame, height, width and channel dimensions, respectively) at each time step $t$ and textual feature $y$ as input. MotionDirector adapts ZeroScope [31] to perform motion customization.
@@
$\min_{\theta_{tp}}\mathbb{E}_{z_{0},y,t,\epsilon}\left[\left\lVert\epsilon-\epsilon_{\theta,\theta_{tp}}(z_{t},y,t)\right\rVert_{2}^{2}\right],$ (1)
@@
where $\epsilon$ denotes the noise added in $z_{t}$ and $\epsilon_{\theta,\theta_{tp}}$ denotes the U-Net $\epsilon_{\theta}$ with temporal LoRAs $\theta_{tp}$. As a result, MotionDirector effectively models the motion concept within the trained temporal LoRAs. At inference, it generates videos simply utilizing the trained temporal LoRAs.
@@
However, we observe MotionDirector tends to overfit the appearance of the reference videos, suggesting that the trained temporal LoRAs still encode the appearance information. It is reasonable as with explicit regularization, the dual-path LoRA adaptation framework cannot guarantee the temporal LoRA does not encode the appearance information. In this paper, we carefully examine the architecture of typical T2V DMs and propose two novel strategies to constrain the temporal LoRA learning, including the temporal attention purification (TAP) and appearance highway (AH).
3.2 Temporal attention purification
@@
In this section, we propose Temporal Attention Purification strategy (TAP) to separate the motion encoding from appearance encoding in adapting temporal transformer block with temporal LoRAs.
@@
Specifically, we assume that in temporal attention module, the pretrained Value embeddings $W_{V}$ are sufficient to serve as the basic components needed for depicting a motion. To generate a new motion, what we need to adapt is the temporal attention, which determines how we combine the Value embeddings in temporal attention module. Thus, in TAP, we choose to utilize temporal LoRAs only to adapt the Query or the Key embeddings to reshape the temporal attention to produce a new motion, without disturbing the Value embeddings. As the temporal attention is only utilized to combine the Value embeddings, we expect reshaping it will not introduce appearance information.
@@
We empirically find that our assumption is reasonable. As shown in Tab. 1, we evaluate the effects of adapting different parts of a temporal transformer block. By adapting different parts of temporal transformer block with LoRAs, we have different generation results. We utilize an off-the-shelf motion classifier [41] to classify the generated videos. Higher accuracy (denoted as “motion” in the table) means higher quality of motions generated to an extent. We also utilize CLIP to evaluate the appearance alignment of generated videos with the text descriptions of users (denoted as “App. Align.”) and those of reference videos (denoted as “App. Leak.”). Higher “App. Align.” and lower “App. Leak.” means better appearance quality of generated videos. From the Tab. 1, we observe that compared to the T2V-DM baseline, adapting Query, Key, Value and the feed-forward network (FFN) of temporal transformer (i.e., MotionDirector) do improves the motion quality, but scarifies the appearance quality very much. Adapting Value and the feed-forward network also harm the appearance modeling. In contrast, reshaping the temporal attention (including reshaping Query, Key or both) yields high-quality videos considering both motion quality and appearance quality. All those results align well with our basic assumption.
@@
Table 1: Impact of different temporal attention modules on generation quality.
@@
Based on our observations, we choose to only adapt the Key embedding of a temporal transformer block with temporal LoRA, which is formulated as
$\min_{\theta_{K}}\mathbb{E}_{z_{0},y,t,\epsilon}\left[\left\lVert\epsilon-\epsilon_{\theta,\theta_{K}}(z_{t},y,t)\right\rVert_{2}^{2}\right].$ (2)
Compared with Eq. 1, the difference is that we only insert temporal LoRA to adapt Key embeddings.
3.3 Appearance Highway
As TAP may not perfectly avoid encoding appearance information in motion LoRAs, we propose using Appearance Highway (AH) strategy to maintain the appearance generation ability of the original T2V-DM and further avoid appearance leakage issues. Specifically, we alter the starting point of each skip connection in spatial-temporal U-Net [8, 14] from the output of each temporal transformers to the output of each spatial transformers.

### v2

@@ Abstract.
Motion customization aims to adapt the diffusion model (DM) to generate videos with the motion specified by a set of video clips with the same motion concept. To realize this goal, the adaptation of DM should be possible to model the specified motion concept, without compromising the ability to generate diverse appearances. Thus, the key to solving this problem lies in how to separate the motion concept from the appearance in the adaptation process of DM. Typical previous works explore different ways to represent and insert a motion concept into large-scale pre-trained text-to-video diffusion models, e.g., learning a motion LoRA, using latent noise residuals, etc. While those methods can encode the motion concept, they also inevitably encode the appearance in reference videos, resulting in weakened appearance generation capability. In this paper, we follow the typical way to learn a motion LoRA to encode the motion concept, but propose two novel strategies to enhance motion-appearance separation, including temporal attention purification (TAP) and appearance highway (AH). Specifically, we assume that in the temporal attention module, the pretrained Value embeddings are sufficient to serve as basic components needed by producing a new motion. Thus, in TAP, we choose only to reshape the temporal attention with motion LoRAs so that Value embeddings can be reorganized to produce a new motion. Further, in AH, we alter the starting point of each skip connection in U-Net from the output of each temporal attention module to the output of each spatial attention module. Extensive experiments demonstrate that compared to previous works, our method can generate videos with appearance more aligned with the text descriptions and motion more consistent with the reference videos. Our homepage can be found in https://github.com/LiuHuijie6410/SeperateMotionFromAppearance.
Generation Models, Text-to-Video Generation, Diffusion models, Motion Customization
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754493isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Information systems Multimedia content creationccs: Computing methodologies Motion captureccs: Computing methodologies Computer vision representationsccs: Computing methodologies Shape representationsccs: Computing methodologies Reconstruction
[…]
Limitations of previous methods.
@@
Figure 1. Limitations of previous methods. Previous methods (Zhao et al., 2023) suffer from the issue of “appearance leakage”, where elements from the reference videos, such as “window”, unexpectedly appear in the generated video.
1. Introduction
@@
Text-to-video diffusion models (Brooks et al., 2024; Blattmann et al., 2023a) have made remarkable progress. With a text input, these models can generate high-quality videos that faithfully reflect the provided prompt. However, relying solely on text prompts, pre-trained text-to-video diffusion models struggle to generate high-fidelity videos for highly customized motion concepts. To address this issue, researchers explored various ways to realize motion customization of text-to-video diffusion models.
@@
Formally, motion customization aims to adapt the diffusion model (DM) to generate videos with the motion specified by a set of video clips (i.e., reference videos) with the same motion concept. To realize this goal, the adaptation of DM should be possible to model the specified motion concept, without compromising the ability to generate diverse appearance. Thus, the key to solving this problem lies in how to separate the motion concept from the appearance in the adaptation process of diffusion models. Typical previous works (Zhao et al., 2023; Ren et al., 2024; Jeong et al., 2024) explore different ways to represent and insert a motion concept into large-scale pre-trained text-to-video diffusion models. For example,  (Zhao et al., 2023) trains motion LoRAs to encode the motion concept from reference videos. Additionally, (Zhao et al., 2023) trains appearance LoRAs to encode the appearance of reference videos. At inference, the appearance LoRAs are abandoned to avoid overfitting to the appearance of reference videos. Other works like (Jeong et al., 2024) utilize residual vectors between consecutive frames to represent a motion. Although these methods can encode the motion concept, there exists no explicit constraint that appearance information should not be learned or encoded into the motion representations. Actually, they inevitably encode the appearance of reference videos (termed as “appearance leakage”), resulting in a weakened appearance generation ability. As shown in Fig. 1 (first row), the background of the reference videos includes some “window”, which has nothing to do with the motion concept. Previous methods (second row), while fitting the motion, also mistakenly fit the appearance of the “window”. As a result, the generated video’s background also includes the “window”, which is not aligned with the text description.
@@
In this paper, we follow the typical way to learn motion LoRAs to encode the motion concept but propose two novel strategies to enhance motion-appearance separation, including temporal attention purification (TAP) and appearance highway (AH). We choose to adopt the typical text-to-video diffusion model (T2V-DM) ZeroScope (Sterling, 2023) or ModelScope (Wang et al., 2023b) to perform motion customization. Specifically, we assume that in the temporal transformers of text-to-video diffusion models, the pre-trained Value embeddings are sufficient to serve as basic components needed to produce a new motion. Thus, in TAP, we choose only to reshape the temporal attention with motion LoRAs so that Value embeddings can be reorganized to produce a new motion. As TAP may not perfectly avoid encoding appearance information in motion LoRAs, we propose AH to maintain the appearance generation ability of original T2V-DM and further avoid appearance leakage issue. Specifically, we alter the starting point of each skip connection in spatial-temporal U-Net (Çiçek et al., 2016; Ho et al., 2022b) from the output of each temporal transformer to the output of each spatial transformers. The underlying logic of AH is that as the skip connection of U-Net mainly conveys high-frequency appearance information (Si et al., 2024), the starting point alternation can provide a “highway” for the hidden states of non-adapted spatial transformers and cut the shortcut from the hidden states of LoRA-adapted temporal transformers. We empirically find that through AH, the hidden states from the decoding branch of spatial-temporal U-Net can encode appearance information more similar to the vanilla T2V-DM and motion information more similar to the T2V-DM adapted with TAP. This means AH largely maintains the appearance generation ability of vanilla T2V-DM, without harming the motion concept learning from reference videos. With the proposed TAP and AH, our method can largely mitigate the appearance leakage issue and generate videos with customized motion and high-fidelity appearance well aligned with the text descriptions, as shown in Fig. 1 (last row). Additionally, we point out that motion modeling occurs in the early steps of the denoising process and propose phased LoRA integration (PIL) to maximize the T2V-DM’s appearance modeling capability. Extensive experiments demonstrate that our method effectively separates motion from appearance and outperforms the previous state-of-the-art.
In a nutshell, our contributions are summarized as follows,
@@
  • We conducted extensive qualitative and quantitative experiments, demonstrating that our method outperforms the previous state-of-the-art.
@@
Figure 2. Overview of our method. This figure illustrates our main contributions: (a) Overall inference process. (b) Temporal Attention Purification where we utilize LoRA only to adapt the Key embeddings to reshape the temporal attention. (c) Appearance Highway where we alter the starting point of skip connection from LoRA-adapted temporal transformer block to non-adapted spatial transformer block.
[…]
Overview of our method.
2. Related Work
2.1. Text-to-Video Generation.
@@ 2. Related Work
Text-to-video generation is a task that involves generating realistic videos based on text prompts. Early research focused on Generative Adversarial Networks (GAN) (Vondrick et al., 2016; Balaji et al., 2019; Tian et al., 2021; Shen et al., 2023), autoregressive transformers(Yan et al., 2021; Hong et al., 2022; Moing et al., 2021; Ge et al., 2022). Recently, diffusion models have emerged as the most popular paradigm (Brooks et al., 2024; An et al., 2023; Esser et al., 2023; Huang et al., 2024a; Zhang et al., 2023c, a, b; Blattmann et al., 2023a). Make-A-Video (Singer et al., 2022) is first trained on labeled images and then on unlabeled videos to address the problem of the lack of paired video-text data. Imagen Video(Ho et al., 2022a) generates high-quality videos using a text-to-video model combined with spatial and temporal super-resolution models. MagicVideo(Zhou et al., 2022) and VideoLDM(Blattmann et al., 2023b) adapt the Latent Diffusion Model’s (Rombach et al., 2022) architecture for video generation. In contrast to these methods, Text2Video-Zero(Khachatryan et al., 2023), AnimateDiff(Guo et al., 2023) leverage the prior knowledge of image generation models to generate videos. Recently, open-source T2V-DM, ModelScope (Wang et al., 2023b) and ZeroScope (Sterling, 2023) have attracted considerable attention.
2.2. Motion Customization Generation.
@@
To generate motion-customized videos, several video editing methods (Ceylan et al., 2023; Yang et al., 2023; Geyer et al., 2023; Qi et al., 2023), including Control-A-Video (Chen et al., 2023) and VideoComposer (Wang et al., 2024), transfer motion from a reference video to new appearances. These approaches primarily modify the appearance without fully capturing the essence of motion, which limits appearance diversity. Tune-A-Video (Wu et al., 2023a) extends text-to-image models into the video domain, enabling customized video generation. VMC (Jeong et al., 2024) uses noise residuals to represent motion information and minimizes residual loss to fine-tune pre-trained text-to-video models. LAMP (Wu et al., 2024) and DreamVideo (Wei et al., 2024) incorporate an additional image paired with editing text as input, serving as guidance for the generation process. However, this approach increases the complexity of user input requirements. To address these challenges, MotionDirector (Zhao et al., 2023) and Customize-a-Video (Ren et al., 2024) introduce prior knowledge of video appearance into their models, encouraging a stronger focus on learning motion rather than appearance.
3. Method
@@
In this paper, we aim to adapt the text-to-video diffusion model (T2V-DM) to perform motion customization, i.e., we need to adapt T2V-DM with reference videos containing one specific motion concept to generate new high-fidelity videos with the same motion concept.
The overview of our method is illustrated in Fig. 2. Our method is based on typical pre-trained T2V-DMs (e.g., ZeroScope (Sterling, 2023) or ModelScope (Wang et al., 2023b)) whose architecture is a 3D U-Net. The basic components for such a 3D U-Net include spatial transformer and temporal transformer, which encode the spatial and temporal information, respectively. Skip connections are utilized to connect the encoders and decoders of the U-Net.
@@
Following MotionDirector, we adopt a dual-path LoRA adaption for training (Sec. 3.1), which tunes the model with low-rank adaptions (LoRAs) while keeping the pre-trained parameters fixed, and finally models the motion information in the LoRAs of temporal transformers. However, such a strategy still inevitably encodes the appearance of the reference videos into temporal LoRAs, resulting in a weakened appearance generation ability. Thus, we propose two effective modifications. First, instead of injecting LoRAs into all linear layers of the attention mechanism in temporal transformers, we propose a Temporal Attention Purification (TAP) strategy (Sec. 3.2). Secondly, at the inference stage, we modify the skip connection between the encoder and decoder of spatial-temporal U-Net with an Appearance Highway strategy (Sec. 3.3), which transfers the hidden states from the spatial transformers rather than the temporal transformers of the encoder to the decoder. Finally, during inference, we adopt a phased LoRA integration trick where we utilize our modified framework to denoise for a few steps and then use the original spatial-temporal U-Net for the latter steps (Sec. 3.4).
3.1. Baseline: Dual-Path LoRA Adaption
@@
In this paper, we adopt MotionDirector (Zhao et al., 2023) as our baseline. MotionDirector utilizes a spatial-temporal U-Net (Çiçek et al., 2016; Ho et al., 2022b), which takes a visual latent code $z_{t}\in\mathcal{R}^{b\times f\times h\times w\times c}$ ($b$, $f$, $h$, $w$, $c$ representing batch size, the number of frame, height, width, and channel dimensions, respectively) at each time step $t$ and textual feature $y$ as input. MotionDirector adapts ZeroScope (Sterling, 2023) to perform motion customization.
@@
(1) $\min_{\theta_{tp}}\mathbb{E}_{z_{0},y,t,\epsilon}\left[\left\lVert\epsilon-\epsilon_{\theta,\theta_{tp}}(z_{t},y,t)\right\rVert_{2}^{2}\right],$
@@
where $\epsilon$ denotes the noise added in $z_{t}$ and $\epsilon_{\theta,\theta_{tp}}$ denotes the U-Net $\epsilon_{\theta}$ with temporal LoRAs $\theta_{tp}$. As a result, MotionDirector models the motion concept within the trained temporal LoRAs. At inference, it generates videos simply utilizing trained temporal LoRAs.
@@
However, we observe that MotionDirector tends to overfit the appearance of the reference videos, suggesting that the trained temporal LoRAs still encode the appearance information. It is reasonable as with explicit regularization, the dual-path LoRA adaptation framework cannot guarantee the temporal LoRA does not encode the appearance information. In this paper, we carefully examine the architecture of typical T2V-DMs and propose two novel strategies to constrain temporal LoRA learning, including temporal attention purification (TAP) and appearance highway (AH).
3.2. Temporal Attention Purification
@@
In this section, we propose Temporal Attention Purification strategy (TAP) to separate motion encoding from appearance encoding in adapting the temporal transformer block with temporal LoRAs.
@@
Specifically, we assume that in the temporal attention module, the pre-trained Value embeddings ($W_{V}$) are sufficient to serve as the basic components needed for depicting a motion. To generate a new motion, what we need to adapt is the temporal attention, which determines how we combine the Value embeddings in the temporal attention module. Thus, in TAP, we choose to utilize temporal LoRAs only to adapt the Query ($W_{Q}$) or the Key embeddings ($W_{K}$) to reshape the temporal attention to produce a new motion, without disturbing the Value embeddings. As the temporal attention is only utilized to combine the Value embeddings, we expect reshaping it will not introduce appearance information.
We probe each component of the temporal transformer block to verify our assumption. Specifically, we evaluate the effects of adapting different parts of a temporal transformer block. By adapting different parts of the temporal transformer with LoRAs, we have different results. We utilize an off-the-shelf motion classifier (Wu et al., 2023c) to classify the generated videos. Higher accuracy (denoted as “motion” in Tab. 1) mean higher quality of the generated motions. We remove the verbs from the edited text prompt and calculate the CLIP scores between the verb-free text prompt and the video as a metric of appearance alignment (App. Align.). We then calculate the CLIP scores between text “a person” and the video as a metric of appearance leakage (App. Leak). For example, we fine-tune the model using several reference videos of a person playing golf, then generate a video with the prompt “a monkey is playing golf on a field full of flowers.” To evaluate appearance alignment, we compute the CLIP similarity between the generated video and the text “a monkey is on a field of flowers”. To measure appearance leakage, we compute the CLIP similarity using the text “a person.”
@@
Higher “App. Align.” and lower “App. Leak.” means better appearance quality of the generated videos. From the Tab. 1, we observe that compared to the T2V-DM, adapting Query, Key, Value and the feed-forward network ($W_{ff}$) of temporal transformer (i.e., MotionDirector) does improve the motion quality, but scarifies the appearance quality. Adapting Value and the feed-forward network also harm the appearance modeling. In contrast, reshaping the temporal attention (including reshaping Query, Key or both) yields high-quality videos considering both motion quality and appearance quality. All those results align well with our assumption.
Table 1. Probing experiments on the temporal attention transformer block. For metric definitions, refer to Sec. 3.2
@@
Based on our observations, we choose to only adapt the Key embedding of a temporal transformer block with temporal LoRAs.
3.3. Appearance Highway